OpenStudy (anonymous):

A box has dimensions of 14 inches long, 1.5 feet wide, and 7 inches high. What is the volume of the box? The formula for the volume is V = l ⋅ w ⋅ h.

OpenStudy (dayakar):

l=14inches w=1.5feet= 1.5 *12 inches (1 feet =12 inches) h= 7 inches v = l*w*h v= 14*1.5*12*7 simplify it